How many grams of N2 will be contained in a 3.0 L flask at STP?


1
Expert's answer
2022-03-13T15:17:14-0400

n = V/22.4

n (N2)=3/22.4=0.13 mol

n = m/M

m(N2)=nxM = 0.13x28=3.64 g